# references.bib

### v1

@@ Abstract
We study the convergence of gradient methods for the training of mean-field single hidden layer neural networks with square loss. Observing this is a separable non-linear least-square problem which is linear w.r.t. the outer layer’s weights, we consider a Variable Projection (VarPro) or two-timescale learning algorithm, thereby eliminating the linear variables and reducing the learning problem to the training of the feature distribution. Whereas most convergence rates for the training of neural networks rely on a neural tangent kernel analysis where features are fixed, we show such a strategy enables provable convergence rates for the sampling of a teacher feature distribution. Precisely, in the limit where the regularization strength vanishes, we show that the dynamic of the feature distribution corresponds to a weighted ultra-fast diffusion equation. Relying on recent results on the asymptotic behavior of such PDEs, we obtain guarantees for the convergence of the trained feature distribution towards the teacher feature distribution in a teacher-student setup.
@@ 1 Introduction
Machine learning methods based on artificial neural networks have recently experienced a significant increase in popularity due to their efficiency in solving numerous supervised or unsupervised learning tasks. This success owes to their capacity to perform feature learning, that is to extract meaningful representations from the data during the training process [Goodfellow-et-al-2016, Chap. 15], standing in contrast with kernel methods for which feature representations are designed by hand and fixed during training [hofmann2008kernel]. Feature learning is believed to play a fundamental role in the generalization performance of neural networks. For example, adaptivity to low-dimensional representations of the data can prevent the curse of dimensionality [bach2017breaking].
However, the process through which features are learned remains largely misunderstood. Indeed, adaptability of the representation comes in neural networks at the price of a nonlinear parameterization, making the training dynamic more difficult to analyze. Specifically, for overparameterized neural networks architecture where the dimension of the parameter space greatly exceeds the number of training samples, recent works have put forward the crucial role played by the choice of scaling w.r.t. the number of parameters in the training dynamic [chizat2019lazy, liu2020linearity, yang2021tensor]. For single hidden layer neural networks, the “kernel regime”, corresponding to a scaling of $1/\sqrt{M}$ where $M$ is the width, has been identified as a scaling for which the model is well-approximated by its linearization around initialization, therefore reducing to a kernel method [jacot_neural_2021]. Relying on the good conditioning of the “Neural Tangent Kernel (NTK)”, this regime provides convergence of gradient descent towards a global minimizer of the risk at a linear rate [allen2019convergence, du2019gradient, lee2019wide, zou_gradient_2020]. However, this regime has also been shown to suffer from a “lazy training” behavior preventing significant modification of the feature distribution and associated to poor generalization guarantees [chizat2019lazy].
@@ 1.1 Mean-field neural networks and two-timescale learning
We consider in this work neural networks with parameter in the parameter space $\Omega$, which is assumed to be either the $n$-dimensional torus $\mathbb{T}^{n}=\mathbb{R}^{n}/\mathbb{Z}^{n}$, or a closed, bounded and convex domain of $\mathbb{R}^{n}$. In the following, we are given a map $\phi:\Omega\times\mathbb{R}^{d}\to\mathbb{R}$, called feature map, and for an integer $M\geq 1$, we define a Single Hidden Layer (SHL) neural network of width $M$ with inner weights $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$ and outer weights $\{u_{i}\}_{1\leq i\leq M}\in\mathbb{R}^{M}$ as the map:
@@
taking inputs in the input space $\mathbb{R}^{d}$ and returning values in the output space $\mathbb{R}$. Thanks to the interchangeability of the indices and to the normalisation factor $1/M$, the above model can be reparameterized in terms of the empirical distribution of the inner weights $\{\omega_{i}\}_{1\leq i\leq M}$. Given an arbitrary probability distribution $\mu\in\mathcal{P}(\Omega)$ on the space of inner weights and a measurable map $u\in L^{1}(\mu)$ we define:
@@
In particular, for the empirical distribution $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$ and the outer weights $\hat{u}(\omega_{i})=u_{i}$ we recover the finite width SHL $F_{\hat{\mu},\hat{u}}=F_{\{(\omega_{i},u_{i})\}}$. Such a “mean-field” model of neural network has been proposed by several authors to study the training of neural network at arbitrary large width [chizat2018global, mei2019mean, rotskoff2019global, sirignano2020mean].
@@
We consider in this work a regression setting where the neural network weights are trained for minimizing the mean square error between the neural network’s output and a target signal. In the following, we are given an input data distribution $\rho\in\mathcal{P}(\mathbb{R}^{d})$, a target signal $Y\in L^{2}(\rho)$ and we define the training risk for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$ as:
$\mathcal{R}(\mu,u)\coloneqq\frac{1}{2}\|F_{\mu,u}-Y\|^{2}_{L^{2}(\rho)}\,.$
@@
Training the neural networks amounts to find parameters $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$ minimizing the training risk $\mathcal{R}(\mu,u)$. Such a task is usually performed using first order optimization methods such gradient descent or stochastic versions of gradient descent on the neural network’s weights [bottou2018optimization].
@@
We study a Variable Projection (VarPro) algorithm where elimination of the linear variable $u$ enables reducing the training of a neural network to the sole problem of learning the feature distribution. Introduced in [golub1973differentiation] for the minimization of separable nonlinear least-squares problems, such a strategy has proven to be efficient in various applications [golub2003separable, osborne2007separable]. The principle of this method can however apply to a broader class of problems [osborne2007separable, aravkin2012estimating] and, for general convex functions, one can show using Schur complement that partial optimization over one variable leads to a better conditioning of the Hessian [calafiore2014optimization, vialard2022partial].
Exploiting the linearity w.r.t. the outer weights in the definition of $F$, it is convenient to read a neural network’s output $F_{\{(\omega_{i},u_{i})\}}(x)=\frac{1}{M}\sum u_{i}\phi(\omega_{i},x)$ as a linear combination of the features $\{\phi(\omega_{i},x)\}_{i=1}^{M}$. From this point of view, neural networks should be compared to kernel methods for which the features are built in advance and fixed during training, whereas only the weights of the linear combination are learned [hofmann2008kernel]. In contrast, both inner weights $\{\omega_{i}\}_{i=1}^{M}$ and outer weights $\{u_{i}\}_{i=1}^{M}$ of a neural network are usually trained. In the following, we refer to the parameters $\omega\in\Omega$ as the neural network’s features and to $\mu\in\mathcal{P}(\Omega)$ as the feature distribution. More generally in the mean-fied limit, for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$, we have:
@@
where we introduced the feature operator $\Phi_{\mu}:u\in L^{1}(\mu)\mapsto\int_{\Omega}u(\omega)\phi(\omega,.)\mathrm{d}\mu(\omega)\in L^{2}(\rho)$. Given a regularization parameter $\lambda>0$, we consider the regularization $\mathcal{R}^{\lambda}$ of the training risk w.r.t. the outer weights defined for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$ as:
$\mathcal{R}^{\lambda}(\mu,u)\coloneqq\frac{1}{2}\|F_{\mu,u}-Y\|^{2}_{L^{2}(\rho)}+\lambda\|u\|^{2}_{L^{2}(\mu)}=\frac{1}{2}\|\Phi_{\mu}\cdot u-Y\|^{2}_{L^{2}(\rho)}+\lambda\|u\|^{2}_{L^{2}(\mu)},$
where we assume $\mathcal{R}^{\lambda}(\mu,u)=+\infty$ if $u\notin L^{2}(\mu)$. One can notice that the problem of minimizing the risk $\mathcal{R}^{\lambda}$ belongs to the class of separable nonlinear least-squares problems. Indeed, given a fixed inner weights’ distribution $\mu\in\mathcal{P}(\Omega)$, the problem of minimizing $\mathcal{R}^{\lambda}$ w.r.t. $u$ is a ridge regression problem and, for $\lambda>0$, there exists a unique solution $u^{\lambda}[\mu]\in\operatorname*{arg\,min}_{u\in L^{2}(\mu)}\mathcal{R}^{\lambda}(\mu,u)$ given by $u^{\lambda}[\mu]\coloneqq(\Phi_{\mu}^{\top}\Phi_{\mu}+2\lambda)^{-1}\Phi_{\mu}^{\top}Y$. Plugging this in $\mathcal{R}^{\lambda}$ gives rise to a reduced risk defined for any $\mu\in\mathcal{P}(\Omega)$ by:
@@
where the infimum is taken to be +$\infty$+ ∞ whenever the signal $Y$ is not in the range of $\Phi_{\mu}$. In the case where $Y\in\operatorname{Range}(\Phi_{\mu})$, this minimization problem admits a unique solution and $\mathcal{L}^{0}(\mu)=\|u^{\dagger}[\mu]\|^{2}_{L^{2}(\mu)}$, where $u^{\dagger}[\mu]=\Phi_{\mu}^{\dagger}Y$ and $\Phi_{\mu}^{\dagger}$ is the generalized pseudo-inverse of $\Phi_{\mu}$ restricted to $L^{2}(\mu)$.
Two-timescale learning
We study the behavior of first order methods such as gradient flow or gradient descent for the minimization of the reduced risk $\mathcal{L}^{\lambda}$ for $\lambda\geq 0$. Whereas the algorithmic complexity of gradient descent might be high as the evaluation of $\mathcal{L}^{\lambda}$ implies the resolution of a linear regression problem, it can also conveniently be be seen as a two-timescale limit of gradient descent performed on the risk $\mathcal{R}^{\lambda}$ [borkar1997stochastic, borkar2008stochastic].
@@
For a neural network of finite width $M\geq 0$ with features $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$, the associated reduced risk is ^$\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}\}_{1\leq i\leq M})\coloneqq\mathcal{L}^{\lambda}(\hat{\mu})$ where $\hat{\mu}$ is the empirical distribution $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$. Then, for an initialization $\{\omega_{i}^{0}\}_{1\leq i\leq M}\in\Omega^{M}$ and a step-size $\eta>0$, the gradient descent algorithm updates the features iteratively following:
$\forall k\geq 0,\forall i\in\{1,...,M\},\quad\omega_{i}^{k+1}=\omega_{i}^{k}-\eta\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}^{k}\}_{1\leq i\leq M})\,.$
@@
For the purpose of theoretical analysis we consider the limit of the gradient descent algorithm when the step-size $\eta$ tends to 00. For an initialization $\{\omega_{i}(0)\}_{1\leq i\leq M}\in\Omega^{M}$ the gradient flow of ^$\hat{\mathcal{L}}^{\lambda}_{f}$ is defined for every time $t\geq 0$ as the solution $\{\omega_{i}(t)\}_{1\leq i\leq M}\in\Omega^{M}$ to the ODE:
@@
$\forall i\in\{1,...,M\},\quad\frac{\mathrm{d}}{\mathrm{d}t}\omega_{i}(t)=-\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}(t)\}_{1\leq i\leq M})\,.$ (6)
Similarly, for a neural network with weights $\{(\omega_{i},u_{i})\}_{1\leq i\leq M}\in(\Omega\times\mathbb{R})^{M}$ the associated risk is ^$\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i},u_{i})\}_{1\leq i\leq M})\coloneqq\mathcal{R}^{\lambda}(\hat{\mu},\hat{u})$, where $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$ and $\hat{u}(\omega_{i})=u_{i}$, and for an initialization $\{(\omega_{i}(0),u_{i}(0))\}_{1\leq i\leq M}$, the gradient flow dynamic reads:
$\forall i\in\{1,...,M\},\quad\left\{\begin{array}[]{cc}\frac{\mathrm{d}}{\mathrm{d}t}\omega_{i}(t)&=-\nabla_{\omega_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i}(t))\}_{1\leq i\leq M})\\[8.0pt] \frac{\mathrm{d}}{\mathrm{d}t}u_{i}(t)&=-\gamma\nabla_{u_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i}(t))\}_{1\leq i\leq M})\end{array}\right.$
where we introduce the timescale parameter $\gamma>0$ controlling the ratio of learning timescales between the parameters $\{\omega_{i}\}_{1\leq i\leq M}$ and $\{u_{i}\}_{1\leq i\leq M}$. When $\gamma<1$ the outer-weights $u_{i}$ are learned “slower” than the features $\omega_{i}$ and reversely, when $\gamma>1$ the outer-weights $u_{i}$ are learned “faster” than the features $\omega_{i}$. In particular, the limiting training dynamic when $\gamma\to+\infty$ corresponds (formally) to the case where the outer weights are learned “instantaneously”, that is, at each time $t\geq 0$, we have $\{u_{i}(t)\}_{1\leq i\leq M}\in\operatorname*{arg\,min}_{u\in\mathbb{R}^{M}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i})\}_{1\leq i\leq M})$. Then by the envelope theorem $\nabla_{\omega_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i}(t))\}_{1\leq i\leq M})=\lambda\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}(t)\}_{1\leq i\leq M})$. The training dynamic eq. 6 can thus be obtained (at least formally) as the training dynamic of a neural network’s features in the limit where $\gamma\to+\infty$. We refer to eq. 6 as the two-timescale regime.
@@
We assume that,
@@
    there exists a measure $\bar{\nu}\in\mathcal{M}(\Omega)$ s.t. $Y=\Phi\star\bar{\nu}$, where $\Phi\star:\mathcal{M}(\Omega)\to L^{2}(\rho)$ is the convolution operator defined by:
    $\forall\nu\in\mathcal{M}(\Omega),\quad\Phi\star\nu\coloneqq\int_{\Omega}\phi(\omega,.)\mathrm{d}\nu(\omega)\;;$ (7)
@@
In such a “teacher-student” framework, we are interested in determining to what extent the teacher feature distribution can be learned by the student neural network. Observe that, under Assumption 1, $\mathcal{L}^{0}$ can be simply expressed in terms of the $\chi^{2}$-divergence between the teacher feature distribution $\bar{\mu}$ and $\mu$. By definition $\chi^{2}(\bar{\mu}|\mu)=\int_{\Omega}\left|\frac{\mathrm{d}\bar{\mu}}{\mathrm{d}\mu}-1\right|^{2}\mathrm{d}\mu$ and:
@@
The Wasserstein gradient flow of $\mathcal{L}^{0}$ then corresponds to nonlinear diffusion equations of the form:

### v2

@@ Abstract
We study the convergence of gradient methods for the training of mean-field single-hidden-layer neural networks with square loss. For this high-dimensional and non-convex optimization problem, most known convergence results are either qualitative or rely on a neural tangent kernel analysis where nonlinear representations of the data are fixed. Using that this problem belongs to the class of separable nonlinear least squares problems, we consider here a Variable Projection (VarPro) or two-timescale learning algorithm, thereby eliminating the linear variables and reducing the learning problem to the training of nonlinear features. In a teacher-student scenario, we show such a strategy enables provable convergence rates for the sampling of a teacher feature distribution. Precisely, in the limit where the regularization strength vanishes, we show that the dynamic of the feature distribution corresponds to a weighted ultra-fast diffusion equation. Recent results on the asymptotic behavior of such PDEs then give quantitative guarantees for the convergence of the learned feature distribution.
@@ 1 Introduction
Machine learning methods based on artificial neural networks have recently experienced a significant increase in popularity due to their efficiency in solving numerous supervised or unsupervised learning tasks. This success owes to their capacity to perform feature learning, that is to extract meaningful representations from the data during the training process [Goodfellow-et-al-2016, Chap. 15], standing in contrast with kernel methods for which feature representations are designed by hand and fixed during training [hofmann2008kernel]. Feature learning is believed to play a fundamental role in the generalization performance of neural networks. For example, adaptivity to low-dimensional representations of the data can prevent the curse of dimensionality [bach2017breaking, ghorbani2020neural].
However, the process through which features are learned remains largely misunderstood. Indeed, adaptivity of the representations comes in neural networks at the price of a nonlinear parameterization, making the training dynamic more difficult to analyze. Specifically, for overparameterized neural network architectures where the dimension of the parameter space greatly exceeds the number of training samples, recent works have put forward the crucial role played by the choice of scaling w.r.t. the number of parameters in the training dynamic [chizat2019lazy, liu2020linearity, yang2021tensor]. For single-hidden-layer neural networks, the “kernel regime”, corresponding to a scaling of $1/\sqrt{M}$ where $M$ is the width, has been identified as a scaling for which the model is well-approximated by its linearization around initialization, therefore reducing to a kernel method [jacot_neural_2021]. Relying on the good conditioning of the “Neural Tangent Kernel (NTK)”, this regime provides convergence of gradient descent towards a global minimizer of the risk at a linear rate [allen2019convergence, du2019gradient, lee2019wide, zou_gradient_2020]. However, this regime has also been shown to suffer from a “lazy training” behavior preventing significant modification of the feature distribution and associated to poor generalization guarantees [chizat2019lazy].
@@ 1.1 Mean-field neural networks and two-timescale learning
We consider in this work neural networks with parameter in the parameter space $\Omega$, which is assumed to be either the $n$-dimensional torus $\mathbb{T}^{n}=\mathbb{R}^{n}/\mathbb{Z}^{n}$, or a closed, bounded and convex domain of $\mathbb{R}^{n}$. In the following, we are given a map $\phi:\Omega\times\mathbb{R}^{d}\to\mathbb{R}$, called feature map, and for an integer $M\geq 1$, we define a single-hidden-layer (SHL) neural network of width $M$ with inner weights $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$ and outer weights $\{u_{i}\}_{1\leq i\leq M}\in\mathbb{R}^{M}$ as the map:
@@
taking inputs in the input space $\mathbb{R}^{d}$ and returning values in the output space $\mathbb{R}$. Thanks to the interchangeability of the indices and the normalisation factor $1/M$, the above model can be reparameterized in terms of the empirical distribution of the inner weights $\{\omega_{i}\}_{1\leq i\leq M}$. Given an arbitrary probability distribution $\mu\in\mathcal{P}(\Omega)$ on the space of inner weights and a measurable map $u\in L^{1}(\mu)$ we define:
@@
In particular, for the empirical distribution $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$ and the outer weights $\hat{u}(\omega_{i})=u_{i}$ we recover the finite width SHL $F_{\hat{\mu},\hat{u}}=F_{\{(\omega_{i},u_{i})\}}$. Such a “mean-field” model of neural network has been proposed by several authors to study the training of neural networks at arbitrary large width [chizat2018global, mei2019mean, rotskoff2019global, sirignano2020mean].
Supervised learning
@@
In the context of supervised learning, training a neural network consists in minimizing a training risk associated to the evaluation of the model on some training data. We consider in this work a univariate regression setting where the neural network weights are trained for minimizing the mean square error with a target signal $Y\in L^{2}(\rho)$ evaluated on training data with distribution $\rho\in\mathcal{P}(\mathbb{R}^{d})$. For a regularization strength $\lambda>0$ and $\mu\in\mathcal{P}(\Omega)$, $u\in L^{1}(\mu)$ we define the training risk as:
$\mathcal{R}^{\lambda}(\mu,u)\coloneqq\frac{1}{2}\|F_{\mu,u}-Y\|^{2}_{L^{2}(\rho)}+\lambda\|u\|^{2}_{L^{2}(\mu)},$ (3)
where we assume $\mathcal{R}^{\lambda}(\mu,u)=+\infty$ if $u\notin L^{2}(\mu)$. Training the neural network then amounts to finding parameters $(\mu,u)\in\operatorname*{arg\,min}\mathcal{R}^{\lambda}$.
Example of applications
Note that the mean-field neural network model of Eq. 2 can be seen as a linear model acting on (signed) measures. Indeed, for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$, we have $F_{\mu,u}=\Phi\star(u\mu)$ where for every finite Borel measure $\nu\in\mathcal{M}(\Omega)$ we define:
$\Phi\star\nu\coloneqq\int_{\Omega}\phi(\omega,.)\mathrm{d}\nu(\omega)\,.$ (4)
In turn, minimization of functionals of the form in Eq. 3 with linear models acting on the space of measures have numerous applications depending on the choice of the feature map $\phi$.
  • Two-layer perceptron: The perceptron model is arguably the prototypical example of a neural network. It consists in considering $\Omega\subset\mathbb{R}^{d+1}$ and a feature map $\phi:(\omega,x)\mapsto\sigma(\omega^{\top}\bar{x})$ where $\bar{x}=(x,1)\in\mathbb{R}^{d+1}$ and $\sigma:\mathbb{R}\to\mathbb{R}$ is some nonlinear activation such as the Rectified Linear Unit (ReLU) or hyperbolic tangent. Owing to their great expressivity [cybenko1989approximation], this class of models is ubiquitous in applications where an unknown signal is to be recovered from data observations.
  • Radial Basis Function (RBF) neural networks and signal deconvolution: RBF neural networks [pereyra2006variable, karamichailidou2024radial] is an example of a simple architecture in which the feature map consists of a translation invariant kernel $k$ i.e. $\Omega\subset\mathbb{R}^{d}$ and $\phi:(\omega,x)\mapsto k(\omega-x)$. The network $F_{\mu,u}$ then implements a convolution with the kernel $k$ and minimization of the risk $\mathcal{R}^{\lambda}$ amounts to solve a form of deconvolution problem. This has important applications in signal processing where one wants to recover an unknown signal given noisy or filtered observations [de2012exact, duval2015exact].
Training with gradient descent and two-timescale learning
@@
In supervised learning, minimization of the training risk is usually performed using first order optimization methods such as gradient descent or stochastic variants on the neural network’s weights [bottou2018optimization].
For a SHL of finite width $M\geq 1$ with weights $\{(\omega_{i},u_{i})\}_{1\leq i\leq M}\in(\Omega\times\mathbb{R})^{M}$ the associated risk is ^$\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i},u_{i})\}_{1\leq i\leq M})\coloneqq\mathcal{R}^{\lambda}(\hat{\mu},\hat{u})$, where $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$ and $\hat{u}(\omega_{i})=u_{i}$. For an initialization $\{(\omega_{i}^{0},u_{i}^{0})\}_{1\leq i\leq M}$, a step-size $\tau>0$ and a timescale parameter $\eta>0$, the gradient descent dynamic reads:
$\forall k\geq 0,\quad\forall i\in\{1,...,M\},\quad\left\{\begin{array}[]{rcl}\omega_{i}^{k+1}&=&\omega_{i}^{k}-M\tau\nabla_{\omega_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}^{k},u_{i}^{k})\}_{1\leq i\leq M})\\[8.0pt] u_{i}^{k+1}&=&u_{i}^{k}-\eta M\tau\nabla_{u_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}^{k},u_{i}^{k})\}_{1\leq i\leq M})\end{array}\right.$ (7)
For the purpose of theoretical analysis we study here the limit of the gradient descent algorithm when the step-size $\tau$ tends to 00. For an initialization $\{(\omega_{i}(0),u_{i}(0))\}_{1\leq i\leq M}$, this gradient flow dynamic reads:
$\forall i\in\{1,...,M\},\quad\left\{\begin{array}[]{rcl}\frac{\mathrm{d}}{\mathrm{d}t}\omega_{i}(t)&=&-M\nabla_{\omega_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i}(t))\}_{1\leq i\leq M})\\[8.0pt] \frac{\mathrm{d}}{\mathrm{d}t}u_{i}(t)&=&-\eta M\nabla_{u_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i}(t))\}_{1\leq i\leq M})\end{array}\right.$ (10)
Note the role of the timescale parameter $\eta>0$ controlling the ratio of learning timescales between inner and outer weights. When $\eta<1$ the outer-weights $u_{i}$ are learned more “slowly” than the inner-weights $\omega_{i}$ and conversely, when $\eta>1$ the outer-weights $u_{i}$ are learned more “quickly” than the inner-weights $\omega_{i}$. In particular, the limiting training dynamics when $\eta\to+\infty$ correspond (formally) to the case where the outer weights are learned “instantaneously”, that is, at each time $t\geq 0$, we have $\{u_{i}(t)\}_{1\leq i\leq M}\in\operatorname*{arg\,min}_{u\in\mathbb{R}^{M}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i})\}_{1\leq i\leq M})$. Such limiting dynamics correspond to a variable projection algorithm.
@@
The Variable Projection (VarPro) algorithm performs elimination of the linear variable $u$ and enables here reducing the training of a neural network to the sole problem of learning the feature distribution. Introduced in [golub1973differentiation] for the minimization of separable nonlinear least squares problems, such a strategy has proven to be efficient in various applications [golub2003separable, osborne2007separable] including the training of neural networks [sjoberg1997separable, pereyra2006variable, newman2021train, karamichailidou2024radial]. A reason for this popularity is that partial optimization over one variable can lead to a better conditioning of the Hessian [sjoberg1997separable, vialard2022partial].
Exploiting here the linearity w.r.t. the outer weights in the definition of $F$, it is convenient to read a neural network’s output $F_{\{(\omega_{i},u_{i})\}}(x)=\frac{1}{M}\sum u_{i}\phi(\omega_{i},x)$ as a linear combination of the features $\{\phi(\omega_{i},x)\}_{i=1}^{M}$. From this point of view, neural networks should be compared to kernel methods for which the features are built in advance and fixed during training, whereas only the weights of the linear combination are learned [hofmann2008kernel]. In contrast, both inner weights $\{\omega_{i}\}_{i=1}^{M}$ and outer weights $\{u_{i}\}_{i=1}^{M}$ of a neural network are usually trained. In the following, we refer to the parameters $\omega\in\Omega$ as the neural network’s features and to $\mu\in\mathcal{P}(\Omega)$ as the feature distribution. More generally in the mean-field limit, for $\mu\in\mathcal{P}(\Omega)$ and $u\in L^{1}(\mu)$, we have:
@@
where we introduced the feature operator $\Phi_{\mu}:u\in L^{1}(\mu)\mapsto\int_{\Omega}u(\omega)\phi(\omega,.)\mathrm{d}\mu(\omega)\in L^{2}(\rho)$. One can thus notice that the problem of minimizing the risk $\mathcal{R}^{\lambda}$ belongs to the class of separable nonlinear least squares problems as, by definition, for a fixed inner weights distribution $\mu\in\mathcal{P}(\Omega)$:
$\mathcal{R}^{\lambda}(\mu,u)=\frac{1}{2}\|\Phi_{\mu}\cdot u-Y\|^{2}_{L^{2}(\rho)}+\lambda\|u\|^{2}_{L^{2}(\mu)}.$
Thus the problem of minimizing $\mathcal{R}^{\lambda}$ w.r.t. $u$ is a ridge regression problem which can be efficiently numerically solved by inverting a linear system. For $\lambda>0$, there exists a unique solution $u^{\lambda}[\mu]\in\operatorname*{arg\,min}_{u\in L^{2}(\mu)}\mathcal{R}^{\lambda}(\mu,u)$ given by $u^{\lambda}[\mu]\coloneqq(\Phi_{\mu}^{\top}\Phi_{\mu}+2\lambda)^{-1}\Phi_{\mu}^{\top}Y$. Plugging this in $\mathcal{R}^{\lambda}$ gives rise to a reduced risk which we define for any $\mu\in\mathcal{P}(\Omega)$ by:
@@
where the infimum is taken to be +$\infty$+ ∞ whenever the signal $Y$ is not in the range of $\Phi_{\mu}$. In the case where $Y\in\operatorname{Range}(\Phi_{\mu})$, this minimization problem admits a unique solution $u^{\dagger}[\mu]=\Phi_{\mu}^{\dagger}\cdot Y$, where $\Phi_{\mu}^{\dagger}$ is the generalized pseudo-inverse of $\Phi_{\mu}$ restricted to $L^{2}(\mu)$, and $\mathcal{L}^{0}(\mu)=\|u^{\dagger}[\mu]\|^{2}_{L^{2}(\mu)}$.
@@
The VarPro algorithm consists here in performing gradient descent over the reduced risk $\mathcal{L}^{\lambda}$. For a neural network of finite width $M\geq 1$ with features $\{\omega_{i}\}_{1\leq i\leq M}\in\Omega^{M}$, the associated reduced risk is ^$\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}\}_{1\leq i\leq M})\coloneqq\mathcal{L}^{\lambda}(\hat{\mu})$, where $\hat{\mu}$ is the empirical distribution $\hat{\mu}=\frac{1}{M}\sum_{i=1}^{M}\delta_{\omega_{i}}$. For an initialization $\{\omega_{i}^{0}\}_{1\leq i\leq M}\in\Omega^{M}$ and a step-size $\tau>0$, the VarPro dynamic reads:
$\forall k\geq 0,\forall i\in\{1,...,M\},\quad\omega_{i}^{k+1}=\omega_{i}^{k}-M\tau\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}^{k}\}_{1\leq i\leq M})\,.$
@@
As before, the gradient flow of ^$\hat{\mathcal{L}}^{\lambda}_{f}$ is the continuous counterpart of gradient descent when the step-size $\tau$ tends to 00. For an initialization $\{\omega_{i}(0)\}_{1\leq i\leq M}\in\Omega^{M}$, it is defined for every time $t\geq 0$ as the solution $\{\omega_{i}(t)\}_{1\leq i\leq M}\in\Omega^{M}$ to the ODE:
@@
$\forall i\in\{1,...,M\},\quad\frac{\mathrm{d}}{\mathrm{d}t}\omega_{i}(t)=-M\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}(t)\}_{1\leq i\leq M})\,.$ (14)
Note that the above gradient can be efficiently calculated numerically once optimization on the outer weights $u_{i}$ has been performed, for example by means of standard automatic differentiation libraries. Indeed, if $\{u_{i}(t)\}_{1\leq i\leq M}\in\operatorname*{arg\,min}_{u\in\mathbb{R}^{M}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i})\}_{1\leq i\leq M})$, then by the envelope theorem $\nabla_{\omega_{i}}\hat{\mathcal{R}}^{\lambda}(\{(\omega_{i}(t),u_{i}(t))\}_{1\leq i\leq M})=\lambda\nabla_{\omega_{i}}\hat{\mathcal{L}}^{\lambda}(\{\omega_{i}(t)\}_{1\leq i\leq M})$. For the same reason, the above dynamic can be seen, at least formally, as the limit of the gradient flow dynamic Eq. 10 over the (unreduced) risk ^$\hat{\mathcal{R}}^{\lambda}$ when the timescale parameter $\eta$ tends to +$\infty$+ ∞. Thus, we equivalently refer to Eq. 14 as the VarPro gradient flow or as the two-timescale regime of gradient flow.
@@
Let $\Phi\star$ be defined by Eq. 4. We assume that,
@@
    there exists a finite measure $\bar{\nu}\in\mathcal{M}(\Omega)$ s.t. $Y=\Phi\star\bar{\nu}$,
@@
In such a “teacher-student” framework, we are interested in determining to what extent the teacher feature distribution can be learned by the student neural network. Observe that, under Assumption 1, $\mathcal{L}^{0}$ can be simply expressed in terms of the $\chi^{2}$-divergence between the teacher feature distribution $\bar{\mu}$ and $\mu$. By definition $\chi^{2}(\bar{\mu}|\mu)=\int_{\Omega}\left|\frac{\mathrm{d}\bar{\mu}}{\mathrm{d}\mu}-1\right|^{2}\mathrm{d}\mu$ and it follows from Eq. 25 that:
@@
The Wasserstein gradient flow of $\mathcal{L}^{0}$ then corresponds to a nonlinear diffusion equation of the form: